A bag contains 4 red, 9 blue and 2 green marbles. Find the probability that a randomly selected marble is blue.
Choose your answer
4/15 3/5 2/15 9/13

Answer by venugopalramana(3286)

A bag contains 4 red, 9 blue and 2 green marbles. Find the probability that a randomly selected marble is blue.
TOTAL =4+9+2=15 MARBLES
BLUE ARE 9 ....SO PROBABILITY =9/15=3/5